OpenStudy (anonymous):

help pwease i have midterm today!! =*( Can someone please explain successive percentages to me. Spot prices for crude oil are rather volatile. From 1998 to 1999, spot prices for crude oil decreased by 28%. From 1999 to 2000, they increased by 106%. What was the percentage change over the two year period from 1998 to 2000?

OpenStudy (phi):

one way: say you start with 100 in 1998 decrease by 28%: 100- 28% of 100= 100-28= 72 1998 100 1999 72 2000 increase by 106% 72+ 1.06*72= 148.32 can you find the total percent change from 1998 to 2000? 100+ 100*x% = 148.32

OpenStudy (anonymous):

oh im so lost... i don't know what step to take next

OpenStudy (anonymous):

is this the formula (New-Old)/Old = % Old + % * Old = New or Old*(1+%)=New Old = New 1 + %

OpenStudy (phi):

use (New-Old)/Old = % (actually this is a fraction. multiply by 100 to get percent) If this topic is a bit fuzzy, http://www.khanacademy.org/math/arithmetic/percents Get some popcorn and watch these.